Determine the answer to the following expression. State your answer in decimal notation.

1,040+400+55.0
Answer #1

In addition and subtraction, number of digits after decimal places in answer must be the same as the lowest number of digit in added numbers.

1040+400+55.0 = 1440+55.0=1495 (since 1040 doesn't contain digits after decimal)